A business begins with $1,000,000 from its owners. It intends to reach a leverage ratio of 3. How much does it expect to borrow?

a) 4,000,000

b)3,000,000

c)2,000,000

d)1,000,000
Answer #1

Answer: Option c is correct
Leverage ratio=Total assets/Equity=(Equity+Debt)/Equity=1+Debt/Equity
Given that the leverage ratio should be 3 and equity=$1000000
Leverage ratio=1+Debt/Equity
3=1+Debt/1000000
=>3-1=Debt/1000000
=>2=Debt/1000000
=>2*1000000=Debt
=>2000000=Debt